How do you compose an equation from a quadratic graph? The vertex is (-1,-9) The y-intercept is -8, The x intercepts are -4 and 2, and the axis of symmetry is -3 on the x axis.
Answer by scott8148(6628)

the general vertex form is ___ y = a(x - h)^2 + k , where (h,k) is the vertex and "a" is a constant term

y = a(x + 1)^2 - 9

-8 = a(0 + 1)^2 - 9 ___ 1 = a(1)^2 ___ 1 = a

y = (x + 1)^2 - 9 ___ or ___ y = x^2 + 2x - 8